# Ayoung Kim, Dept. of Mechanical Engineering, SNU, Seoul, S. Korea

### v3

@@ 4.2 Sequences
MOANA dataset encompasses two distinct regional environments: South Korea and Singapore. The data collected in South Korea is from Ulsan, an area characterized by predominantly industrial structures. The port area is developed for cargo transport, and the surroundings are densely packed with factory sites and artificial structures. This provided a highly structured environment, which we refer to as the Port sequence.
In contrast, the dataset from Singapore is collected in the Harborfront area, representing a largely unstructured environment. Apart from the berthing point, the surroundings are dominated by natural elements such as trees and rocks, making localization algorithms more challenging. Given the varying sizes of islands in this region, we have designated this sequence as the Island sequence. A detailed explanation of each sequence and the associated challenges is provided below.
4.2.1 Port Sequence:
The Port sequences represent a highly structured, industrial environment characterized by strong and consistent detections of the surrounding area. This sequence is primarily designed for odometry or SLAM applications, where generating a reliable map of the port is the main objective. It is the most straightforward sequence for algorithm testing, with the primary challenge being the mitigation of wave-induced wobble, which leads to inconsistencies in radar data. As radar systems rely on reflected signals, this wobble can disrupt the RCS (RCS) continuity of sensor measurements.
Additionally, the presence of numerous anchored large vessels introduces both opportunities and challenges for tracking. The radar-reflective coatings on these ships make them ideal objects for tracking in open water. However, proximity to these vessels can lead to significant multipath effects, complicating the tracking process.
(i) Near Port: In the Near Port sequence, the dataset primarily emphasizes W-band radar due to its superior performance in short-range detection, as X-band radar exhibits limitations in this condition. W-band radar is, therefore, the predominant sensor for capturing the surrounding environment in these sequences. A significant challenge arises at the halfway point, where severe multipath effects generate ghost objects, a critical obstacle to achieving robust navigation.
@@ 4.2.1 Port Sequence:
Figure 6: Ground truth trajectories for all sequences. The Port sequences (b, c) were collected in South Korea, while the Seaside and Island sequences (d, e, f, g, h) were acquired in Singapore. All sequences share instant common locations except for the Island Expedition sequence (h), which encompasses three sub-sequences (e, f, g). Trajectories commence in red and terminate in blue.
@@ 4.2.2 Island Sequence:
The Island sequences present an unstructured, sporadic environment dominated by natural elements. Due to these features, surrounding detections are inconsistent, making it challenging to recognize the same locations reliably. The dataset includes three distinct island environments: single island, twins island, and complex island. Each sequence contains overlapping areas with others, facilitating global place recognition across the dataset. In addition to island sequences, the sequence also includes seaside areas, providing a mix of both unstructured natural environments and more structured coastal regions.
(i) Single Island: The Single Island sequence features the shortest route in the island dataset, consisting of a simple loop around the island. The yacht completes two laps before returning to the berthing point.
(ii) Twins Island: Twins island is located at the farthest distance and experiences sporadic data loss in the W-band radar. However, as the vessel navigates through the narrow waterway between the islands, W-band radar becomes essential for robust positioning and detection, compensating for the limitation of the X-band radar.
@@
(iv) Seaside: Lastly, the Seaside sequence incorporates both beachside environments and a container loading zone. This sequence is particularly useful for analyzing the impact of both structured and unstructured environments within a single dataset.
Table 3: Absolute Trajectory Error for W-band and X-band radars. Due to challenging data in other sequences, only the Near Port sequence was evaluated using both W-band and X-band radar odometry algorithms. The W-band radar demonstrates lower error in the Near Port sequence, attributable to its superior close-range detection accuracy compared to the X-band radar. In contrast, the X-band radar exhibits robust performance across all other sequences not covered by the W-band radar sensor.
@@ 5.1 W-band radar odometry (CFEAR)
For the benchmark test of W-band radar odometry, we employed the CFEAR (Adolfsson et al., 2022) method, a state-of-the-art approach for W-band radar odometry estimation. Although our goal was to evaluate the entire trajectory, empty-object data in certain areas made this infeasible. As a result, the W-band radar odometry had to be limited to sections where the vessel remained near the shoreline. The results are depicted in Figure. 7 and Table. 3.
@@ 7 Conclusion
MOANA presents the first maritime multi-radar dataset incorporating scanning radars of different bandwidths. Our dataset facilitates the use of W-band radar in oceanic environments while maintaining vessel ego-motion tracking by leveraging existing X-band radar. Addressing the limitations of W-band radar, W-band radar also mitigates the vulnerability of X-band radar during berthing operations. Our benchmark result also demonstrated that integrating the X-band and W-band radar can further enhance the performance of maritime navigation tasks such as berthing, sailing, and docking. Seven sequences in this dataset are expected to support advancements in autonomous navigation systems. In future updates, we aim to extend the dataset with sequences incorporating temporal variance to evaluate robustness under time-differentiated conditions.
@@
This work was supported by the Technology Innovation Program (1415187329, Development of autonomous driving connectivity technology based on sensor-infrastructure cooperation) funded By the Ministry of Trade, Industry & Energy (MOTIE, Korea), and the research in the Singapore region was supported by Defence Science and Technology Agency (DSTA)

### v4

@@ 4.2 Sequences
Figure 7: Ground truth trajectories for all sequences. The Port sequences (b, c) were collected in South Korea, while the Seaside and Island sequences (d, e, f, g, h) were acquired in Singapore. All sequences share slight common locations except for the Island Expedition sequence (h), which encompasses three sub-sequences (e, f, g). Trajectories commence in red and terminate in blue.
The MOANA dataset encompasses two distinct regional environments: South Korea and Singapore. The data collected in South Korea is from Ulsan, an area characterized by predominantly industrial structures. The port area is developed for cargo transport, and the surroundings are densely packed with factory sites and artificial structures. This provided a highly structured environment, which we refer to as the Port sequence.
In contrast, the dataset from Singapore is collected in the Harborfront area, representing a largely unstructured environment. Apart from the berthing point, the surroundings are dominated by natural elements such as trees and rocks, making localization algorithms more challenging. Given the varying sizes of islands in this region, we have designated this sequence as the Island sequence. The overall trajectory for each sequence is depicted in Figure. 7, with detailed explanations of each sequence and the associated challenges provided below.
4.2.1 Port Sequence:
The Port sequences represent a highly structured, industrial environment characterized by strong and consistent detections of the surrounding area. This sequence is primarily designed for odometry or SLAM (SLAM) applications, where generating a reliable map of the port is the main objective. The GNSS data exhibits periodic delays due to the sensor interference and regional characteristics in the Port sequences, highlighting the critical importance of achieving accurate sensor-driven SLAM results.
These sequences are the most straightforward for algorithm testing, with the primary challenge being the mitigation of wave-induced wobble, which leads to inconsistencies in radar data. As radar systems rely on reflected signals, this wobble can disrupt the RCS (RCS) continuity of sensor measurements. Additionally, the presence of numerous anchored large vessels introduces both opportunities and challenges for tracking. The radar-reflective coatings on these ships make them ideal objects for tracking in open water. However, proximity to these vessels can lead to significant multipath effects, complicating the tracking process.
(i) Near Port: In the Near Port sequence, the dataset primarily emphasizes W-band radar due to its superior performance in short-range detection, as X-band radar exhibits limitations in this condition. W-band radar is, therefore, the predominant sensor for capturing the surrounding environment in these sequences. A significant challenge arises at the midpoint of the entire route, where severe multipath effects generate ghost objects, a critical obstacle to achieving robust navigation.
@@ 4.2.2 Island Sequence:
The Island sequences depict an unstructured environment characterized by sparse, irregular radar returns predominantly arising from natural elements. Due to these features, surrounding detections are inconsistent, making it challenging to recognize the same locations reliably. The dataset includes three distinct island environments: single island, twins island, and complex island. Each sequence contains overlapping areas with others, facilitating global place recognition across the dataset. In addition to island areas, the sequence also includes seaside areas, providing a mix of both unstructured natural environments and more structured coastal regions.
(i) Single Island: The Single Island sequence features the shortest route in the island dataset, consisting of a simple loop around the island collected at night. The yacht completes two laps before returning to the berthing point.
(ii) Twins Island: Twins island is located at the farthest distance and experiences sporadic data loss in the W-band radar. However, as the vessel navigates through the narrow waterway between the islands, W-band radar becomes essential for robust positioning and detection, compensating for the limitation of the X-band radar. Twins island sequence begins in darkness and gradually transitions to brightness by the end.
@@
(iv) Seaside: Lastly, the Seaside sequence incorporates both beachside environments and a container loading zone. This sequence is particularly useful for analyzing the impact of both structured and unstructured environments within a single dataset. Seaside sequence starts in bright daylight and ends in darkness.
Table 3: Absolute Trajectory Error for W-band and X-band radars. Due to challenging data in other sequences, only the Near Port sequence was evaluated using both W-band and X-band radar odometry algorithms. The W-band radar demonstrates lower error in the Near Port sequence, attributable to its superior short-range detection accuracy compared to the X-band radar. In contrast, the X-band radar exhibits robust performance across all other sequences not covered by the W-band radar sensor.
@@ 5.1 W-band radar odometry (CFEAR)
For the benchmark test of W-band radar odometry, we employed the CFEAR (Adolfsson et al., 2022) method, a state-of-the-art approach for W-band radar odometry estimation. Although our goal was to evaluate the entire trajectory, lacking distinct objects in certain regions made this infeasible. As a result, the W-band radar odometry had to be limited to sections where the vessel remained near the shoreline. The results are depicted in Figure. 8 and Table. 3.
@@ 7 Conclusion
MOANA presents the first maritime multi-radar dataset incorporating scanning radars of different bandwidths. Our dataset facilitates the use of W-band radar in oceanic environments while maintaining vessel ego-motion tracking by leveraging existing X-band radar. The high-resolution imaging provided by the W-band radar appears to mitigate the perceptual limitations of the X-band radar during operations requiring precise sensing capabilities, such as berthing or docking. Our benchmark results indicate that while W-band radar delivers enhanced odometry estimation during near harbor sequences, its performance in wide-area scenarios is limited. Thus, a hybrid approach integrating X-band and W-band radar data could potentially improve maritime navigation tasks in a complementary manner. Seven sequences in this dataset are expected to support advancements in autonomous navigation systems. In future updates, we aim to extend the dataset with sequences incorporating temporal variance to evaluate robustness under time-differentiated conditions.
@@
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00461409), and the research in the Singapore region was supported by Defence Science and Technology Agency (DSTA)